OpenStudy (anonymous):

For [-12,12] f(x)=x^5(x+3)^4 on which interval is the function increasing?

OpenStudy (anonymous):

f(x)=x^5(x+3)^4 f'(x) = 5x^4 (x+3)^4 + 4(x+3)^3 x^5 =x^4 (x+3)^3 (5(x+3) +4x) =x^4 (x+3)^3 (9x +15) =3(x^4)(x+3)^3 (3x+5) to check if f(x) is increasing or decreasing, we have to find out the turning points first. How to find it? Since turning points have slope =0, set f'(x)=0 3(x^4)(x+3)^3 (3x+5) =0 x=0, -3 or -5/3 then, we can get ranges : -3<x<-5/3 or -5/3<x<0 or x>0 for -3<x<-5/3, put any values in this range to f'(x). For example, we sub. x=-2 into f'(x).Then, we can get f'(x)<0. Similarly,for -5/3<x<0 , f'(x) >0 for x>0, f'(x) >0 Note that f'(x)>0 means that f(x) is strictly increasing f'(x)<0 means that f(x) is strictly decreasing Then we now can figure out that for -5/3<x<0 and 0<x<12, f(x) is increasing

OpenStudy (anonymous):

THANKS